OpenStudy (anonymous):

Solve 4y - y + 11 = 38. 9 -9 3.8 8.

OpenStudy (anonymous):

y =9

OpenStudy (anonymous):

A

OpenStudy (anonymous):

How did you get that? I've been staring at it for like ever. What do you do when you have two variables?

OpenStudy (anonymous):

4y - y + 11 = 38 3y + 11 = 38 (subtract 11 from both sides) 3y = 38 - 11 3y = 27 (divide both sides by 3) y = 9

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Oh okay I get it, you subtracted the y from the 4y okay thanks!!

OpenStudy (anonymous):

thats what we are here for :)

OpenStudy (anonymous):

always simplify your problem by combining like terms

OpenStudy (♪chibiterasu):

@Jelia97 Don't rush when helping people, it's better to go slower and show them the steps on how to solve.